OpenStudy (anonymous):

The perimeter of rectangle is 80 inches...the area is 396 square inches...find length and width

OpenStudy (lgbasallote):

perimeter of rectangle = 2L + 2W = 80 area = LW = 396 you can use substitution to solve...need guide?

OpenStudy (anonymous):

i think i can get it

OpenStudy (lgbasallote):

just tell me when you need help :)

OpenStudy (anonymous):

i dont understand what you plugin for L & W

OpenStudy (anonymous):

2L + 2W = 80 LW = 396 for LW = 396 L=396/W put this into 2L + 2W = 80 2(396/W) + 2W =80 396/W + W =40 396 +W^2 = 40W W^2 -40W -396 =0 solve for W, reject -ve value of W Then sub the value of W into L=396/W to find L Is the above what you mean, @lgba

OpenStudy (anonymous):

thanks coco...thats it!